OpenStudy (anonymous):

Use the functions to answer the question. f(x) = 10 – x What is the value of (f ◦ f)(3)? 1 3 7 8

OpenStudy (shinalcantara):

you just have to substitute the value of x for the given function f(x) = 10 - x where x=3

OpenStudy (anonymous):

Oh, awesome! So 7. Thanks!!! <3

OpenStudy (shinalcantara):

i think you're not yet done since it implies composite function. therefore you need to plug 7 to the same equation f(x) = 10 - x where x this time is 7

OpenStudy (anonymous):

Oh, I see!! Thank you. I always forget that step. So now it is 3! lol thanks again. :)

OpenStudy (shinalcantara):

yw :)